OpenStudy (anonymous):

mrs.siimms told her class to find twoo consecutive even integers such that twice the lesser of two integers is 4 less than two times the grater integer write and solve and equation to find the integers and does the equation have on solution, no solutions, or is it and identity? Explain

OpenStudy (anonymous):

it has infinite solutin..

OpenStudy (anonymous):

every next even no is 2 greater than the previous..so when u make it twice then the gifference is always 4

OpenStudy (anonymous):

lat the smaller is x and larger=x+2 2(x)=2(x+2)-4 2x=2x it holds for every x so infinite solution